You are thrown from a motorcycle and skid in a straight line to a complete stop on a rough

gravel path. A measurement of the bloody skidmark reveals that it is 11 m long. What is

the average force did the gravel exert on your body? Assume that your mass is 55.0 Kg

and your initial speed was 30 km/h.


1
Expert's answer
2022-06-01T12:49:12-0400
"F=ma=\frac{mv^2}{2d}\\
=\frac{55.0*8.33^2}{2*11}=174\:\rm N"